find the laplace transform of y"-4y'+9y=t, y(0)=0,y'(0)=1
1
Expert's answer
2020-05-26T18:03:28-0400

The given equation is y'' - 4y' + 9y = t, and y(0) = 0 and y'(0) = 1.


We know, L[t] = 1/s²

And, L[(dn/dtn)y(t)] = sn L[y(t)] - sn-1 y(0) - sn-2 y'(0) - ..... - yn-1(0)


Calculating L[y"] :

L[y"] = s² L[y] - s¹ y(0) - s⁰ y'(0)

i.e., L[y"] = s² L[y] - 0 - 1

L[y"] = s² L[y] - 1


Calculating L[y'] :

L[y'] = s¹ L[y] - s⁰ y(0)

L[y'] = s L[y]


Rewriting the given equation as,

y" - 4y' + 9y - t = 0

Calculating the laplace transform :

L[ y" - 4y' + 9y - t] = L[y"] - 4 L[y'] + 9 L[y] - L[t] = 0

(s² L[y] - 1) - 4(sL[y]) + 9 L[y] - (1/s²) = 0

(s² - 4s + 9)L[y] - 1 - 1/s² = 0

(s² - 4s + 9)L[y] = 1 + 1/s²

(s² - 4s + 9)L[y] = (s² + 1)/s²

L[y] = (s² + 1)/(s²(s² - 4s + 9))

"\therefore" L[y] = 1/(s² - 4s + 9) + 1/(s²(s² - 4s + 9))